OpenStudy (anonymous):

Solve the square root of x+31=x+1

OpenStudy (anonymous):

do you know how to get rid of a square root sign?

OpenStudy (anonymous):

no

OpenStudy (saifoo.khan):

\[\sqrt{x+31} = x+1\]Like that?

OpenStudy (anonymous):

Yes

OpenStudy (anonymous):

you square each side so you have \[(\sqrt{x+31})^{2} = (x+1)^{2}\]

OpenStudy (anonymous):

now what happens when you square a square root?

OpenStudy (anonymous):

It goes away?

OpenStudy (anonymous):

exactly, so it just becomes x+31 = \[(x+1)^{2}\]

OpenStudy (anonymous):

now you don't have a square root so you should be able to multipy out your quadratic and solve for x as normal.

OpenStudy (anonymous):

Ok. Thanks! :)

OpenStudy (anonymous):

no problem! so don't forget, if you want to get rid of a square root just square both sides!

OpenStudy (anonymous):

So do I multiply (x+1)(x+1)?